1. Return on total assets A company reports the following income statement and balance sheet information...

1.

Return on total assets

A company reports the following income statement and balance sheet information for the current year:

Net income $172,560
Interest expense 30,450
Average total assets 2,010,000

Determine the return on total assets. If required, round the percentage to one decimal place.
_______%

2.

Common Stockholders' Profitability Analysis

A company reports the following:

Net income $190,000
Preferred dividends 7,600
Average stockholders' equity 1,407,407
Average common stockholders' equity 852,336

Determine (a) the the return on stockholders’ equity and (b) the return on common stockholders’ equity. If required, round your percentages to one decimal place.

a. The rate earned on stockholders' equity ____%
b. The rate earned on common stockholders' equity ____%

3.

Earnings per Share and Price-Earnings Ratio

A company reports the following:

Net income $307,000
Preferred dividends $17,000
Shares of common stock outstanding 20,000
Market price per share of common stock $194.3

a. Determine the company's earnings per share on common stock. Round your answer to the nearest cent. Use the rounded answer of requirement a for subsequent requirement, if required.
$________

b. Determine the company's price-earnings ratio. Round to one decimal place.

_________
Answer #1

1) Return on total assets = (net income+Interest expense)/Average total assets = (172560+30450)/2010000 = 10.1%

2) The rate earned on stockholders' equity = Net income/Average stockholder's equity = 190000/1407407 = 13.5%

Average common stockholders' equity = (190000-7600)/852336 = 21.4%

3) earning per share of common Stock = (307000-17000)/20000 = 14.50

Price earning ratio = 194.3/14.5 = 13.4 times